Please help solve the equation, and check the solution for
3/8x - 1/5x +10= 3/10x

Answer by stanbon(75887)

3/8x - 1/5x +10= 3/10x
------
Multiply thru by 40 to get:
15x - 8x +400 = 12x
5x = 400
x = 80
--------------------
Cheers,
Stan H.